Graph:
y^2/25-x^2/4=1

Answer by lwsshak3(11628)

Graph:
y^2/25-x^2/4=1
This is an equation of a hyperbola with vertical transverse axis. y-term leads
Its standard form: (y-k)^2/a^2-(x-h)^2/b^2=1, (h,k)=(x,y) coordinates of center
For given equation:
center: (0,0)
a^2=25
a=5
vertices: (0,0±a)=(0,0±5)=(0,-5) and (0,5)
..
b^2=4
b=2
length of conjugate axis=2b=4
see graph below:
y=±(25+25x^2/4)^.5